Project

General

Profile

NDOS CommissionMtng 20130108

Attending:Xuebing, Steve Magill, Jaraslov, Kurt, Jon, Jan, Susan, Nick Peter, Vipin

News:
-MC Decommissioning
Xuebing: Reports no progress

APD Test is isolated from APD test stand

-RC transition
-Long holiday runs
-Power script and reboot
-New firmware

UPDATES:
-DataCheck Update (http://nusoft.fnal.gov/nova/datacheck/febshutoff/2012-12-28_2013-01-07.txt)
Thursday
14393 -
14394 - 12 day
14399 - Peter's harmless "test" of LED seemed to have cause 4 feb drops and empty events
8258 - lost 220 channels

NDOS processes were hosed on next run

Jan found script time inconsistencies. Using UTC in the HW DB. Can the DB enforce the time zones?

2-1 46 shutoff in multiple runs (Someone should check it out)

Firmware test on Thursday failed: DSO scans hung with new DCM firmware alone. The changes were rolled back.

-Installation Plans for the week

-DAQ Update and Plans for the week
DAQ shakedown - (Pending)

-DCS Update and Plans for the week

APDTestStand: Leon says APDs on the Teststand will remain in the near term.

-Jon Paley mentions updating the shifter information

Reboot-dcm-all needs to get rid of the old mc pieces (Ron rebuilt the host list to avoid conflicts with APDTest)

PowerUtils added to testRelforOperations on NDOS master. power-off/on-ndos scripts now available. Experts are advised to use DCS gui for power control and avoid command line control.

AOB
Jan
Remote control room? Jan talked about test remote control. Jarek is handling this -> Peter will sending documentation. Shift trail.